OpenStudy (anonymous):

3.) -2/3(c-12)=??? 7)2x(3y+-6)=??? 10.)5x-3x+2xy+31x+-18xy=?? 5.)3/4(12x+16) 7.)(16+13x)10 2.)8(3m+4n 8.)(14+7a) 1/7

OpenStudy (anonymous):

Whats the question?

OpenStudy (anonymous):

those are the questions...

OpenStudy (anonymous):

oh.rewrite into simplest terms using distributive property???????

OpenStudy (anonymous):

Haha betterrrr...Ok, one second.

OpenStudy (anonymous):

if you're helping pleaseee go in order of how I posted them.... thnx(:

OpenStudy (anonymous):

Ill do each one starting from the top: \[-\frac{2}{3}(c-12)=-\frac{2}{3}*c-\frac{2}{3}(-12)=-\frac{2c}{3}+8=8-\frac{2c}{3}\] The second has "\(\pm\)". Is that intentional?

OpenStudy (anonymous):

+- is intentional.. on the one u did jw if the answer is 8-2c/3????

OpenStudy (anonymous):

Yeah?

OpenStudy (anonymous):

ok thanks, imworking out your problem onn my worksheet..

OpenStudy (anonymous):

\[2x(3y+-6)=2x(3y-6)=(2*3*x*y-2*6*y)=6xy-12y\]

OpenStudy (anonymous):

wait please

OpenStudy (anonymous):

on #3. 2c/3=8 cant u multiply the 3 times 8 to simplify that further because I think it can go down further?????

OpenStudy (anonymous):

Nooo, because you had the expression \(-\frac{2}{3}(c-12)\). There was no equals sign in the first place. Where did that one come from haha

OpenStudy (anonymous):

on one of the past lessons I think it was order of operations pemdas the teacher said we can always multiply denominator iwasnt sure if these problems were the same my baddddd... I got 3 and 7 now.. pleaseeee help w/10

OpenStudy (anonymous):

Alright!

OpenStudy (anonymous):

10.\[\eqalign{&\phantom{=.}5x-3x+2xy+31x-18xy \\&=2x+2xy+31x-18xy \\ &=2x+31x+2xy-18xy\\ &=33x-16xy}\] ...not much simplifying to do

OpenStudy (anonymous):

Tell me when your ready for next

OpenStudy (anonymous):

ok ready.. THANKS soooo much for help (:

OpenStudy (anonymous):

No problem :) happy to help: \[\eqalign{& \phantom{=.}\frac{3}{4}(12x+16) \\ &=9x+12 \\ }\]

OpenStudy (anonymous):

Sorry im answering a lot of questions at the same time lol

OpenStudy (anonymous):

7 pleaseeee

OpenStudy (anonymous):

is #7 multiplied by 10 or is it to the tenth power ?

OpenStudy (anonymous):

\[10(13x+16)=130x+160\]

OpenStudy (anonymous):

No, its a question on distributive property. Multiplication. That would be a really pain that one eh?

OpenStudy (anonymous):

I got 130x+160 because 10(16)=160 am I right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

please help w/2 and 8.. and then I have 5more similar questions pleaseeee stay on here though):

OpenStudy (anonymous):

Alright!

OpenStudy (anonymous):

\[8(3m+4n)=24m+32n\]

OpenStudy (anonymous):

ok I understood that one.. please help w/8 ad then 5 more

OpenStudy (anonymous):

\[\frac{1}{7}(14+7a)=2+a\]

OpenStudy (anonymous):

23.)5xy(-2x)(-3y) 27.)-15a(4b)(2bc) 24.)7a(-4ab)(2a) 26.) -cd(5c)(-7cd) 28.) 7x(-9xy)(-8y)

OpenStudy (anonymous):

your hes veryyyyyyy much appreciated.. (:

OpenStudy (anonymous):

\[5xy(-2x)(-3y)=30x^2y^2\]

OpenStudy (anonymous):

\[-15a(4b)(2bc)=-120ab^2c\]

OpenStudy (anonymous):

\[7a(-4ab)2a=-56a^3b\]

OpenStudy (anonymous):

\[-(cd)(5c)(-7cd)=35c^2d^2\]

OpenStudy (anonymous):

\[7x(-9xy)(-8y)=504x^2y^2\]

OpenStudy (anonymous):

DONE! haha I got to go.